# RIME: Robust Preference-based Reinforcement Learning with Noisy Preferences

### v2

@@ 4.2 Warm Start
Sample selection methods frequently suffer from accumulated error due to incorrect selection, which underscores the need of good initialization for the discriminator to effectively differentiate between samples at initial. Meanwhile, we observe a marked degradation of performance during transition from pre-training to online training (see Figure 3). This gap is clearly observed under noisy feedback settings and is fatal to robustness. It is exacerbated when following the most widely-adopted backbone algorithm, PEBBLE, which resets the Q-network and only retains the pre-trained policy after the pre-training phase. The issue arises because the biased reward model, trained on noisy preferences, biasedly optimize the Q-network through minimizing Bellman residual. This, in turn, offers a poor learning signal for the policy, erasing any gains made during pre-training.
Figure 3: Performance degradation during transition on Walker-walk (left) and Quadruped-walk (right) with 30% noisy preferences. We pre-train an agent using SAC for 20k steps. Warm start method shows smaller transition gap and faster recovery.
Inspired by these observations, we propose to warm start the reward model to facilitate a smoother transition from pre-training to online training. Specifically, we pre-train the reward model to approximate intrinsic rewards during pre-training phase. Because the output layer of the reward model typically uses the tanh activation function (Lee et al., 2021b), we firstly normalize the intrinsic reward to the range $(-1,1)$ as follows:
@@
where $0<\delta\ll 1$. $\bar{r}$ and $\sigma_{r}$ are mean and standard deviation of the intrinsic rewards, respectively. Then the agent receives the reward $r^{\text{int}}_{\text{norm}}$ and stores each tuple $(\mathbf{s}_{t},\mathbf{a}_{t},r^{\text{int}}_{\text{norm}},\mathbf{s}_{t+1})$ in a replay buffer, denoted as $\mathcal{D}_{\text{pretrain}}$. During the reward model update, we sample batches of $(\mathbf{s}_{t},\mathbf{a}_{t})$ along with all encountered states $\mathcal{S}=\{\mathbf{s}|\mathbf{s}\text{ in }\mathcal{D}_{\text{pretrain}}\}$ for nearest neighbor searches. The loss function for updating the reward model $\hat{r}_{\psi}$ is given by the mean squared error as:
@@
Attributed to warm start, both the Q-network and reward model are aligned with intrinsic rewards, allowing for the retention of all knowledge gained during pre-training (i.e.policy, Q-network, and reward model) for subsequent online training. Moreover, the warm-started reward model contains more information than random initialization, enhancing the discriminator’s ability at initial.
@@ 5 Experiments
Figure 4: Learning curves on locomotion tasks from DMControl, where each row represents a specific task and each column corresponds to a different error rate $\epsilon$ setting. SAC serves as a performance upper bound, using a ground-truth reward function unavailable in PbRL settings. The corresponding number of feedback in total and per session are show in Table 1. The solid line and shaded regions respectively denote mean and standard deviation of episode return, across ten runs.
@@ 5.1 Setups
Implementation Details. For the hyperparameters of RIME, we fix $\alpha=0.5$, $\beta_{\min}=1$ and $\beta_{\max}=3$ in the lower bound $\tau_{\text{lower}}$, and fix the upper bound $\tau_{\text{upper}}=3\ln(10)$ for all experiments. The decay rate $k$ in $\tau_{\text{upper}}$ is $1/30$ for DMControl tasks, and $1/300$ for Meta-world tasks, respectively. Other hyperparameters are kept the same with PEBBLE. For the sampling of queries, we use the disagreement sampling scheme for all PbRL algorithms, following the setting in Christiano et al. (2017). For the implementation of baselines, we use their corresponding publicly released repositories (see Table 6 for source codes). The feedback amount in total and per query session in each environment with specified error rate are detailed in Table 1.
For each task, we run all algorithms independently for ten times and report the average performance along with the standard deviation. Tasks from Meta-world are measured on success rate, while tasks from DMControl are measured on ground-truth episode return. More details on the algorithm implementation are provided in Appendix D.2.
Table 1: Feedback amount in each environment with specified error rate. Value column refers to feedback amount in total / per session.
@@ 5.2 Results
Table 2: Results on tasks from Meta-world and DMControl with “mistake” teacher. The result shows the mean and standard deviation of metric (i.e.episode return for DMControl tasks and success rate for Meta-world tasks) across all five error rates within 10 runs.
@@
For robotic manipulation tasks, we consider three tasks from Meta-world: Button-press, Sweep-into, and Hammer. For locomotion tasks, we choose three environments from DMControl: Walker-walk, Cheetah-run, and Quadrupted-walk. Figure 2 and Figure 4 show the learning curves of RIME and baselines on Meta-world and DMControl tasks with five error rates, respectively. Table 2 shows the mean and standard deviation of metrics across the five error rates.
Since some preferences are corrupted, we observe that there is a gap between all PbRL methods and the best performance (i.e.SAC with task reward), but RIME exceeds the PbRL baselines by a large margin in almost all environments. Especially, RIME remains effective in cases where all baselines struggle, such as Button-press with $\epsilon=0.2$, Hammer with $\epsilon=0.3$ and Walker with $\epsilon=0.3$, etc. These results demonstrate that RIME significantly improves robustness against noisy preferences. We also observe that although some feedback-efficient baselines based on PEBBLE perform comparable or even exceed PEBBLE in low-level noise, they become ineffective as error rates rise. Additionally, Table 2 shows that PEBBLE is a robust algorithm second only to RIME. This reveals that the pursuit of feedback efficiency leads to over-reliance on feedback quality.
@@ 5.3 Ablation Study
Table 3: Results on tasks from Meta-world and DMControl with 4 other types of (noisy) teacher. The result shows the mean and standard deviation of metric (i.e.episode return for DMControl tasks and success rate for Meta-world tasks) averaged over 5 runs.
@@
Performance with more types of (noisy) teachers. To investigate whether our method can generalize to more situations, we evaluate RIME, PEBBLE and MRN with other four types of teachers proposed by Lee et al. (2021a): Oracle, Skip, Equal, and Myopic. “Oracle” teacher provides ground-truth preferences. “Skip” teacher will skip the query if the cumulative rewards of segments are small. “Equal” teacher will give equal preference $\tilde{y}=(0.5,0.5)$ if the difference between the cumulative rewards of two segments are small. “Myopic” teacher focuses more on the behavior at the end of segments. More details of these four teachers are shown in Appendix D.3. We report mean and standard deviation across five runs in Table 3. We found that RIME not only perform the best when teachers can provide ambiguous or wrong labels (Equal and Myopic), it is also comparable with baselines on correct labels (Oracle and Skip). Based on the superior performance of RIME with multiple teachers, it has stronger chances of performing well with real teachers as well (Lee et al., 2021a).
@@
Comparison with other robust training methods. Since the reward learning in PbRL is posed as a classification problem, we migrate the robust training methods in Machine Learning to compare with RIME. We consider a sample selection method: adaptive denoising training (ADT) (Wang et al., 2021), two robust loss functions: Mean Absolute Error (MAE) (Ghosh et al., 2017) and t-CE (Feng et al., 2021), and a robust regularization method: label smoothing (LS) (Wei et al., 2021), as our baselines. ADT drops a-$\tau(t)$ proportion of samples with the largest cross-entropy loss at each training iteration, where $\tau(t)=\min(\gamma t,\tau_{\max})$. We set $\tau_{\max}=0.3$, $\gamma=0.003$ and $0.0003$ for tasks from DMControl and Meta-world, respectively. MAE loss if formulated as $\mathcal{L}^{\mathtt{MAE}}(\psi)=\operatorname*{\mathbb{E}}[|\tilde{y}-P_{\psi}|]$, while t-CE loss is formulated as $\mathcal{L}^{\mathtt{t}\text{-}\mathtt{CE}}(\psi)=\operatorname*{\mathbb{E}}[\sum_{i=1}^{t}\frac{(1-\tilde{y}^{\top}\cdot P_{\psi})^{i}}{i}]$. Label smoothing method replace $\tilde{y}$ in Equation (2) with $(1-r)\cdot\tilde{y}+\frac{r}{2}\cdot[1,1]^{\top}$. We adopt $t=4$ for t-CE loss and $r=0.1$ for label smoothing respectively. All baselines are implemented based on PEBBLE. Table 4 shows the result on four tasks with “Mistake” teacher and error rate as $\epsilon=0.3$. We observe that label smoothing is almost fails in this setting. Sample selection methods (RIME and ADT) work better compared to other types of methods, and RIME still outperform baselines. The reason is that adding tolerance for out-of-distribution data, i.e.$\tau_{\text{unc}}$ in Equation (5), is more adaptable to RL’s training process.
@@
Table 4: Results of different robust training methods on tasks from Meta-world and DMControl with “mistake” teacher and error rate as $\epsilon=0.3$. The result shows the mean and standard deviation of metric averaged over 5 runs.
@@
Performance with real non-expert human teachers. The ultimate goal of improving robustness is to make PbRL more suitable for humans. To investigate this, following Christiano et al. (2017); Lee et al. (2021b), we conduct a group of experiments with real non-expert human teachers on Hopper to do backflip. In this experiments, we invite five students in unrelated majors, who are blank to the robotic tasks, to perform online annotation. We only tell them the objective of the task (i.e.teach the agent to do backflip) with nothing else and no further guidance. We utilize their annotations to train RIME and PEBBLE. See Appendix D.4 for more details on the annotation protocol. The feedback amount in total and per session are 100 and 10, respectively.
We employ a hand-crafted reward function designed by experts (Christiano et al., 2017) as the ground-truth scripted teacher. We find that compared to ground-truth preferences, the error rate of our non-expert annotations reached nearly 40%. The learning curves are shown in Fig. 5a. We find that RIME significantly outperforms PEBBLE when learning from actual non-expert human teachers and successfully performs consecutive backflips using 100 non-expert feedback, as shown in Figure 5b.
@@
Component analysis. We perform ablation study to individually evaluate each technique in RIME: warm start (WS), lower bound $\tau_{\text{lower}}$, and upper bound $\tau_{\text{upper}}$ of KL divergence. We present results in Table 5 in which we compare the performance of removing each component from RIME. We observe that warm start is crucial for robustness when the number of feedback is quite limited (i.e.on Walker-walk). This is because the limited samples restrict the capability of the reward model, leading to more rounds of queries to cross the transition gap. Moreover, it might be hard to distinguish for discriminator at initial with limited samples, which urges for good initialized reward models.
The lower bound $\tau_{\text{lower}}$ for filtering trustworthy samples is important in high error rate ($\epsilon=0.3$ on Walker-walk and Button-press) and adequate feedback (on Button-press) situations. The upper bound $\tau_{\text{upper}}$ for flipping labels always brings some improvements in our ablation experiments. The full algorithm outperforms every other combination in most tasks. Additionally, the results show that although the contribution of warm start and denoising discriminator vary in different environments with low-level noise, they are both effective and their combination proves essential for the overall success of our method in environments with high-level noise.
@@ 6 Conclusion
In this paper, we present RIME, a robust algorithm for preference-based reinforcement learning (PbRL) designed for effective reward learning from noisy preferences. Unlike previous research which primarily aims to enhance feedback efficiency, RIME focuses on improving robustness by employing a sample selection-based discriminator to dynamically denoise preferences. To reduce accumulated error due to incorrect selection, we utilize a warm-start method for the reward model, enhancing the initial capability of the denoising discriminator. The warm-start approach also serves to bridge the performance gap during transition, facilitating a seamless transition from pre-training to online training. Our experiments show that RIME substantially boosts the robustness of the state-of-the-art PbRL method across a range of complex robotic manipulation and locomotion tasks. Ablation studies further demonstrate that the warm-start approach is crucial for both robustness and feedback efficiency. We believe that RIME has the potential to broaden the applicability of PbRL by leveraging preferences from non-expert users or crowd-sourcing platforms.

### v3

@@ 4.2 Warm Start
Sample selection methods usually suffer from accumulated errors due to incorrect selection, which highlights the need for good initialization. Meanwhile, we observe a significant degradation in performance during the transition from pre-training to online training (see Figure 2). This gap is clearly observable under noisy feedback settings and is fatal to robustness. It is exacerbated when following the most widely-adopted backbone, PEBBLE, which resets the Q-network and only retains the pre-trained policy after the pre-training phase. Because the Q-network is optimized with a biased reward model trained on noisy preferences to minimize the Bellman residual, this biased Q-function leads to a poor learning signal for the policy, erasing gains made during pre-training.
Inspired by these observations, we propose to warm start the reward model to facilitate a smoother transition from pre-training to online training. Specifically, we pre-train the reward model to approximate intrinsic rewards during the pre-training phase. Because the output layer of the reward model typically uses the tanh activation function (Lee et al., 2021b), we firstly normalize the intrinsic reward to the range $(-1,1)$ as follows:
@@
where $0<\delta\ll 1$. $\bar{r}$ and $\sigma_{r}$ are the mean and standard deviation of the intrinsic rewards, respectively. Then the agent receives the reward $r^{\text{int}}_{\text{norm}}$ and stores each tuple $(\mathbf{s}_{t},\mathbf{a}_{t},r^{\text{int}}_{\text{norm}},\mathbf{s}_{t+1})$ in a replay buffer, denoted as $\mathcal{D}_{\text{pretrain}}$. During the reward model update, we sample batches of $(\mathbf{s}_{t},\mathbf{a}_{t})$ along with all encountered states $\mathcal{S}=\{\mathbf{s}|\mathbf{s}\text{ in }\mathcal{D}_{\text{pretrain}}\}$ for nearest neighbor searches. The loss function for updating the reward model $\hat{r}_{\psi}$ is given by the mean squared error as:
@@
Thanks to warm start, both the Q-network and reward model are aligned with intrinsic rewards, allowing for the retention of all knowledge gained during pre-training (i.e.,policy, Q-network, and reward model) for subsequent online training. Moreover, the warm-started reward model contains more information than random initialization, enhancing the discriminator’s ability initially.
Figure 2: Performance degradation during transition on Walker-walk (left) and Quadruped-walk (right) with 30% noisy preferences. We pre-train an agent using SAC for 20k steps. The warm start method shows a smaller transition gap and faster recovery.
@@ 5 Experiments
Figure 3: Learning curves for robotic manipulation tasks from Meta-world, where each row represents a specific task and each column corresponds to a different error rate $\epsilon$. SAC serves as a performance upper bound, using a ground-truth reward function unavailable in PbRL settings. The corresponding number of feedback in total and per session are shown in Table 1. The solid line and shaded regions respectively denote the mean and standard deviation of the success rate, across ten runs.
(a) $\epsilon=0.1$
(b) $\epsilon=0.15$
(c) $\epsilon=0.2$
(d) $\epsilon=0.25$
(e) $\epsilon=0.3$
Figure 4: Learning curves on locomotion tasks from DMControl, where each row represents a specific task and each column corresponds to a different error rate $\epsilon$ setting. SAC serves as a performance upper bound, using a ground-truth reward function unavailable in PbRL settings. The corresponding number of feedback in total and per session are shown in Table 1. The solid line and shaded regions respectively denote the mean and standard deviation of episode return, across ten runs.
@@ 5.1 Setups
Implementation Details. For the hyperparameters of RIME, we fix $\alpha=0.5$, $\beta_{\min}=1$ and $\beta_{\max}=3$ in the lower bound $\tau_{\text{lower}}$, and fix the upper bound $\tau_{\text{upper}}=3\ln(10)$ for all experiments. The decay rate $k$ in $\tau_{\text{upper}}$ is $1/30$ for DMControl tasks, and $1/300$ for Meta-world tasks, respectively. Other hyperparameters are kept the same as PEBBLE. For the sampling of queries, we use the disagreement sampling scheme for all PbRL algorithms, following the setting in Christiano et al. (2017). For the implementation of baselines, we use their corresponding publicly released repositories (see Table 8 for source codes). The feedback amount in total and per query session in each environment with specified error rate are detailed in Table 1.
For each task, we run all algorithms independently ten times and report the average performance along with the standard deviation. Tasks from Meta-world are measured on success rate, while tasks from DMControl are measured on ground-truth episode return. More details on the algorithm implementation are provided in Appendix D.2.
Table 1: Feedback amount in each environment with specified error rate. The “value” column refers to the feedback amount in total / per session.
@@ 5.2 Results
Table 2: Results on tasks from Meta-world and DMControl with “mistake” teacher. The result shows the mean and standard deviation of metric (i.e.,episode return for DMControl tasks and success rate for Meta-world tasks) across all five error rates within 10 runs.
@@
For robotic manipulation tasks, we consider three tasks from Meta-world: Button-press, Sweep-into, and Hammer. For locomotion tasks, we choose three environments from DMControl: Walker-walk, Cheetah-run, and Quadruped-walk. Figure 3 and Figure 4 show the learning curves of RIME and baselines on Meta-world and DMControl tasks with five error rates, respectively. Table 2 shows the mean and standard deviation of metrics across the five error rates.
Since some preferences are corrupted, we observe that there is a gap between all PbRL methods and the best performance (i.e.,SAC with task reward), but RIME exceeds the PbRL baselines by a large margin in almost all environments. Especially, RIME remains effective in cases where all baselines struggle, such as Button-press with $\epsilon=0.2$, Hammer with $\epsilon=0.3$, and Walker with $\epsilon=0.3$, etc. These results demonstrate that RIME significantly improves robustness against noisy preferences. We also observe that although some feedback-efficient baselines based on PEBBLE perform comparable to or even exceed PEBBLE in low-level noise, they become ineffective as the error rate rises. Additionally, Table 2 shows that PEBBLE is a robust algorithm second only to RIME. These results reveal that the pursuit of feedback efficiency leads to over-reliance on feedback quality.
@@ 5.3 Ablation Study
Table 3: Results on tasks from Meta-world and DMControl with 4 other types of (noisy) teacher. The result shows the mean and standard deviation of metric (i.e.,episode return for DMControl tasks and success rate for Meta-world tasks) averaged over 5 runs.
@@
Performance with more types of (noisy) teachers. To investigate whether our method can generalize to more situations, we evaluate RIME, PEBBLE, and MRN with the other four types of teachers proposed by Lee et al. (2021a): Oracle, Skip, Equal, and Myopic. “Oracle” teacher provides ground-truth preferences. “Skip” teacher will skip the query if the cumulative rewards of segments are small. “Equal” teacher will give equal preference $\tilde{y}=(0.5,0.5)$ if the difference between the cumulative rewards of two segments is small. “Myopic” teacher focuses more on the behavior at the end of segments. More details of these four teachers are shown in Appendix D.3. We report mean and standard deviation across five runs in Table 3. We found that RIME not only performs the best when teachers can provide ambiguous or wrong labels (Equal and Myopic), but it is also comparable with baselines on correct labels (Oracle and Skip). Based on the superior performance of RIME with multiple teachers, it has better chances of performing well with real teachers as well (Lee et al., 2021a).
@@
Comparison with other robust training methods. Since the reward learning in PbRL is posed as a classification problem, the robust training methods in Machine Learning could be migrated to compare with RIME. Consider a sample selection method: adaptive denoising training (ADT) (Wang et al., 2021), two robust loss functions: Mean Absolute Error (MAE) (Ghosh et al., 2017) and t-CE (Feng et al., 2021), and a robust regularization method: label smoothing (LS) (Wei et al., 2021), as our baselines. ADT drops a-$\tau(t)$ proportion of samples with the largest cross-entropy loss at each training iteration, where $\tau(t)=\min(\gamma t,\tau_{\max})$. We set $\tau_{\max}=0.3$, $\gamma=0.003$, and $0.0003$ for tasks from DMControl and Meta-world, respectively. MAE loss if formulated as $\mathcal{L}^{\mathtt{MAE}}(\psi)=\operatorname*{\mathbb{E}}[|\tilde{y}-P_{\psi}|]$, while t-CE loss is formulated as $\mathcal{L}^{\mathtt{t}\text{-}\mathtt{CE}}(\psi)=\operatorname*{\mathbb{E}}[\sum_{i=1}^{t}\frac{(1-\tilde{y}^{\top}\cdot P_{\psi})^{i}}{i}]$. Label smoothing method replace $\tilde{y}$ in Equation (2) with $(1-r)\cdot\tilde{y}+\frac{r}{2}\cdot[1,1]^{\top}$. We adopt $t=4$ for t-CE loss and $r=0.1$ for label smoothing respectively. All baselines are implemented based on PEBBLE. Table 4 shows the result on four tasks with “Mistake” teacher and error rate as $\epsilon=0.3$. Additional experiments with fixed lower bound $\tau_{\text{lower}}$ are provided in Appendix E. We observe that label smoothing almost fails to handle corrupted labels in our experiments. Sample selection methods (RIME and ADT) work better compared to other types of methods, and RIME still outperforms baselines. The reason is that the dynamic threshold with tolerance for out-of-distribution data is particularly well-suited to the RL training process.
@@
Table 4: Results of different robust training methods on tasks from Meta-world and DMControl with “mistake” teacher and error rate as $\epsilon=0.3$. The result shows the mean and standard deviation of the metric averaged over 5 runs.
@@
Performance with real non-expert human teachers. The ultimate goal of improving robustness in PbRL is to better align with human users. To investigate how RIME performs with non-expert humans, we conduct experiments on Hopper utilizing non-expert human instructors, following the approach of Christiano et al. (2017); Lee et al. (2021b). Specifically, we selected five students with no prior knowledge on robotics from unrelated majors to provide annotations in an online setting. These students were instructed solely on the objective: to train an agent to perform backflips and received no additional information or guidance on the task. Their annotations were later used to train the algorithms RIME and PEBBLE. The feedback amount in total and per session are 100 and 10 respectively. For further details on the annotation protocol, refer to Appendix D.4.
We employ a hand-crafted reward function designed by experts (Christiano et al., 2017) as the ground-truth scripted teacher. We notice that compared to ground-truth preferences, the error rate of our non-expert annotations reached nearly 40%. The learning curves are shown in Fig. 5a. We find that RIME significantly outperforms PEBBLE when learning from actual non-expert human teachers and successfully performs consecutive backflips using 100 non-expert feedback, as shown in Figure 5b.
@@
Trade-off between sample efficiency and robustness. To quantify the trade-off between sample efficiency and robustness, we conduct experiments with RIME, where we either held the feedback volume constant and increase the error rate, or maintain the error rate while increasing the feedback volume, as shown in Tables 5 and 6, respectively. Table 5 indicates a progressive decline in performance with rising error rates, with a notable deterioration at an error rate of 0.3 across all four environments. Table 6 demonstrates that increasing the feedback volume improves performance, particularly at an error rate of 0.3, where doubling the feedback approximately doubles the performance gains. The same analysis repeated for PEBBLE is shown in Appendix E.
Table 5: Results of RIME as the error rate increases with constant amount of feedback, across 5 runs.
Environment Feedback Error rate
volume 0.1 0.2 0.3
Walker 500 909
$\pm$ 132
806
$\pm$ 162
493
$\pm$ 91
Quadruped 2000 484
$\pm$ 166
400
$\pm$ 166
117
$\pm$ 24
Button-press 10000 99.9
$\pm$ 0.3
92.2
$\pm$ 15.6
38.8
$\pm$ 33.8
Hammer 20000 46.5
$\pm$ 43.9
41.2
$\pm$ 48.0
1.8
$\pm$ 3.1
Table 6: Results of RIME as the feedback volume increases with constant error rate, across 5 runs. $N$ refers to the minimal feedback volume for each environment shown in Table 1.
Domain Environment Error rate Feedback volume
[…]
DMControl Walker 0.2 806
$\pm$ 162
894
$\pm$ 80
0.3 493
$\pm$ 91
741
$\pm$ 139
Quadruped 0.2 400
$\pm$ 166
477
$\pm$ 152
0.3 117
$\pm$ 24
301
$\pm$ 184
Meta-world Button-press 0.2 92.2
$\pm$ 15.6
93.4
$\pm$ 13.6
0.3 38.8
$\pm$ 33.8
80.0
$\pm$ 27.7
Component analysis. We perform an ablation study to individually evaluate each technique in RIME: warm start (WS), lower bound $\tau_{\text{lower}}$, and upper bound $\tau_{\text{upper}}$ of KL divergence. We present results in Table 7 in which we compare the performance of removing each component from RIME. We observe that warm start is crucial for robustness when the number of feedback is quite limited (i.e.,on Walker-walk). This is because the limited samples restrict the capability of the reward model, leading to more rounds of queries to cross the transition gap. Moreover, identifying whether the sample is corrupted or not is challenging for the discriminator initially due to the limited training data, underscoring the need for well-initialized reward models.
The lower bound $\tau_{\text{lower}}$ for filtering trustworthy samples is important in high error rate ($\epsilon=0.3$ on Walker-walk and Button-press) and adequate feedback (on Button-press) situations. The upper bound $\tau_{\text{upper}}$ for flipping labels always brings some improvements in our ablation experiments. The full algorithm outperforms every other combination in most tasks. Additionally, the results show that although the contribution of warm start and denoising discriminator vary in different environments with low-level noise, they are both effective and their combination proves essential for the overall success of RIME in environments with high-level noise.
@@ 6 Conclusion
In this paper, we present RIME, a robust algorithm for preference-based reinforcement learning (PbRL) designed for effective reward learning from noisy preferences. Unlike previous research which primarily aims to enhance feedback efficiency, RIME focuses on improving robustness by employing a sample selection-based discriminator to dynamically denoise preferences. To reduce accumulated error due to incorrect selection, we utilize a warm-start method for the reward model, enhancing the initial capability of the denoising discriminator. The warm-start approach also facilitates a seamless transition from pre-training to online training. Our experiments show that RIME substantially boosts the robustness of the state-of-the-art PbRL method across a wide range of complex robotic manipulation and locomotion tasks. Ablation studies further demonstrate that the warm-start approach is crucial for both robustness and feedback efficiency. We believe that RIME has the potential to broaden the applicability of PbRL by leveraging preferences from non-expert users or crowd-sourcing platforms.
Limitations. The intrinsic challenge of PbRL with noisy preferences is the trade-off between sample-efficiency and robustness. As shown in Table 1, 5, and 6, RIME still needs to increase the amount of feedback in total to perform reasonably as the error rate increases. Therefore, exploring how to achieve better trade-offs is an interesting future direction. Another limitation is that RIME introduces several hyperparameters in the denoising discriminator, resulting in extra tuning efforts for optimal performance. Moreover, the divergence between the noise in actual human preferences and the noise models proposed by BPref deserves further investigation. While attaining high performance across multiple simulated teachers suggests a likelihood of good performance under real human teachers, it remains imperative to investigate simulated noise models that more closely align with the characteristics of real-world human preference noise.
Impact Statement
Compared to natural language processing, control tasks typically demand higher-quality human feedback (Kim et al., 2022). Our work reduces the difficulty of annotating human preferences for control tasks, allowing for the presence of noise in preferences and thereby alleviating the requirement of domain knowledge for annotators. This enables human preferences for control tasks to be sourced from crowd-sourcing platforms or ordinary users, potentially expanding the application scope of PbRL.
Acknowledgments
We are grateful to the anonymous reviewers and editors for their insightful suggestions. This work was partially supported by the National Natural Science Foundation of China under Grants 62271485, U1909204 and 62303462, the Provincial Key Research and Development Program of Zhejiang (project number: 2022C01129), Ningbo International Science and Technology Cooperation Project (2023H020), and Beijing Natural Science Foundation under Grant L233005. Lastly, special thanks go to Tongtian Yue, Ruixi Qiao, and Yueyang Yang for their assistance in data annotation, engaging discussions, and meticulous review of draft versions.